# Massless Quasiparticles in Bogoliubov-de Gennes Systems

### v1

@@ II.4 Conditions for the existence of Zero modes
Now we consider BdG systems on the lattice having a symmetry group $G_{f}$, and assume that a fixed number of complex fermion bases (namely local Wannier orbitals) are placed in certain Wyckoff positions in the unit cell. According to App.C, for any given momentum $\bm{k}$, the rep of $F_{f}(k)$ can be read out from the generalized band representation, and then we can know whether there exist zero modes at the $\bm{k}$. If zero modes do exist, the degeneracy must be even because the dimension of the Hamiltonian $\mathscr{H}_{k}$ is even and the nonzero energy levels appear in pairs $\{\varepsilon,-\varepsilon\}$. In the following theorem 2 (see App.F for proof), however, we show that if the $G_{f}(k)$ is ‘diagonal’ in the particle-hole sector, only minimal nonzero modes are stable at a given momentum $\bm{k}$.
@@
For any BdG system whose symmetry group $G_{f}(k)$ is ‘diagonal’ in the particle-hole sector, all zero modes at the momentum point $\bm{k}$ can be adiabatically removed without breaking any symmetry.
Hence, for non-interacting SCs/SFs whose symmetry groups are magnetic space groups, zero modes at high symmetry points/lines are not symmetry-protected. However, as shown in App.F.3, (effective) charge conjugation symmetry can change the energy spectrum. In App.G, we prove that at a given $\bm{k}$ having $\tilde{\mathcal{P}}$ symmetry, there always exists a (effective) charge-conjugation symmetry $\mathcal{C}$ that turns certain set of minimal nonzero modes into a set of zero modes, called the $\mathcal{C}$-minimal zero modes. This conclusion is summarized in the following theorem 3.
@@
At a given $\tilde{\mathcal{P}}$-symmetric momentum $\boldsymbol{k}$ (i.e. $\tilde{\mathcal{P}}\in F_{f}(k)$) of an arbitrary BdG system, any minimal nonzero modes (except for one special case) located at the momentum $\bm{k}$ can be turned into zero modes by adding a single charge-conjugation symmetry to the original BdG system. The exception is when $\tilde{\mathcal{P}}$ is anti-unitary while $G_{f}(k)$ is unitary, and the irrep $D(F_{f}(k))$ carried by a set of irreducible nonzero modes is of $\mathbb{H}$ class. In that case, one needs two charge-conjugation operations anti-commuting with each other to gain zero modes.
The charge conjugation is the simplest charge operation that is non-diagonal in the particle-hole sector. The existence of zero modes in the above theorem can be generalized to other situations where the charge conjugation operation $\mathcal{C}$ is replaced by general non-diagonal charge operations. For instance, such charge operation can be an element $\mathcal{D}\in SU(2)_{c}$ with $\mathcal{D}^{3}=E$.
On the other hand, zero modes can also appear at unfixed momentum. For instance, a level crossing of irreducible nonzero modes carrying two nonequivalent reps of $G_{f}(k)$ can give rise to a set of zero modes, as stated in the following theorem 4.
@@
On a high symmetry line of the BZ having a unitary symmetry $\mathcal{M}$, if the eigenspaces $\{\varepsilon,-\varepsilon\}$ of a minimal nonzero modes carry different quantum numbers (i.e. characters of irreps) of $\mathcal{M}$, then the level crossing of $\varepsilon,-\varepsilon$ gives rise to a set of stable zero modes whose momentum is not fixed.
@@
(1) presence of (effective) charge-conjugation symmetry or symmetries such that the Hilbert space contains a set of irreducible zero modes or sets of uncoupled irreducible zero modes at high symmetry point $\boldsymbol{k}$.
@@
(3) occurrence of $\pi$-quantized Berry phase protected by ($\mathcal{I}^{\prime}\mathcal{T}^{\prime}$) symmetry with $(\mathcal{I}^{\prime}\mathcal{T}^{\prime})^{2}=1$. Here $\mathcal{I}^{\prime}$ and $\mathcal{T}^{\prime}$ respectively stand for the inversion operation and the time-reversal operation dressed by spin or/and charge operations.
The condition (1) can be verified using the band representation theory for SCs/SFs/QSLs provided in App.C. The last two conditions are illustrated via concrete lattice models in Sec.IV. Especially, the $\mathcal{I}^{\prime}\mathcal{T}^{\prime}$ symmetry enforced $\pi$-quantized Berry phase can give rise to nodal-point like quasiparticles in 2-dimensions and nodal-line structures in 3-dimensions. This is a topological origin of the zero modes (usually linear dispersive), and is closely related to the zero modes at general $\bm{k}$ points enforced by symmetry indicators Po et al. [2017], Bradlyn et al. [2017a], Ono and Watanabe [2018], Geier et al. [2020], Ono et al. [2021], Ono and Shiozaki [2022]. The (projective) representation theory of symmetry groups discussed in the present work provides a different and complementary mechanism to obtain zero modes and gapless quasiparticles.
@@ III.1 $k\cdot p$ Hamiltonian
The above procedure can be applied to the 0-th order $k\cdot p$ terms, i.e., to judge if an irrep gives rise to a set of zero modes and nonzero modes. This can be done by replacing the rep $({\rm v}\times{\rm{\overline{ph}}})$ in eq.(III.1) with the identity rep or the particle-hole rep, namely, if $a_{I}=1,a_{\rm ph}=0$, then the rep corresponds to a set of irreducible zero modes; if if $a_{I}=1,a_{\rm ph}\neq 0$, then the rep corresponds to a set of irreducible nonzero modes, and the obtained $\Lambda$ matrices are independent symmetric perturbation Hamiltonians.
@@
Table 2: Linear Reps of the lattice rotation ($l$ stands for angular momentum), time reversal $\mathcal{T}^{\prime}$, spatial inversion $\mathcal{I}^{\prime}$, charge conjugation $\mathcal{C}$ and particle-hole transformation $\mathcal{P}$ carried by probe fields: chemical potential $\lambda$, momentum $\boldsymbol{k}$, electric field $\boldsymbol{E}$, magnetic field $\boldsymbol{B}$, quadratic dispersion $k_{\alpha}k_{\beta}$, strain field $\epsilon_{\alpha\beta}$, and so on. The last column is the ‘particle-hole rep’ of operations in $F_{f}(k)$.
Prob fields $\lambda$ $\boldsymbol{E}$ $\boldsymbol{B}$ $\boldsymbol{k}$ $k_{\alpha}k_{\beta}$ $\epsilon_{\alpha\beta}$ $D^{\rm{(ph)}}$
@@ III.2 Physical properties
The above method can also be applied to study the physical properties of the quasiparticles, namely, the response to external fields $\boldsymbol{f}$, such as electric field $\boldsymbol{E}$, magnetic field $\boldsymbol{B}$, strain field $\epsilon_{\alpha\beta}$, and so on. As summarized in Tab.2, the prob fields carry linear reps of $g\in F_{f}(k)$, namely
@@
Similar to (18), one has $D(g)K_{s(g)}\Lambda_{\alpha}K_{s(g)}D(g)^{-1}=\sum_{\beta}D^{{\rm(ph)}\times(\mu)}_{\beta\alpha}(g)\Lambda_{\beta}$. The matrices $\Lambda_{\alpha}$ can be obtained using the method introduced in the previous subsection.
Due to the ‘particle-hole symmetry’ $\mathcal{P}$ and the effective charge conjugation symmetry $\mathcal{C}$, the response of the Bogoliubov quasiparticles to external field is different from quasiparticles in semimetals. For instance, the chemical potential $\lambda$ is a scaler field under space-time operations, but it does not carry an identity rep of the group $F_{f}^{c}(k)$, hence the response matrix $\Lambda_{0}$ is not simply an identity matrix. If $\mathcal{P}$ is represented as $\tau_{x}\otimes I$, then the response matrix $\Lambda_{0}$ for $\lambda$ takes the form $\tau_{z}\otimes I$. Hence changing the chemical potential may either shift the position of the nodal point or fully gap out the quasiparticle.
Generally, the external fields may give mass to the gapless quasiparticles. The resulting gapped state may have nontrivial topological properties, such as nontrivial Chern numbers. These information are maintained in $k\cdot p$ model and the response matrix mentioned above. In the following section, we will illustrate all of the above results by several concrete lattice models.
@@ IV Lattice Models: gapless quasiparticles in BdG systems
In this section, we present four models in two-dimensional lattice systems to illustrate the mechanisms and methods proposed in the previous sections. The first three models discuss zero modes protected by an effective charge conjugation $\mathcal{C}$, while the fourth model is a mean field Hamiltonian of $Z_{2}$ QSLs in which the projective symmetry group contains symmetry elements that are non-diagonal in the particle-hole sector.
@@ IV.1 Square lattice: $G_{b}=p4gm\times Z_{2}^{\mathcal{T}}$
The first model is a spin-1/2 SC/SF of the CI class. We consider the square lattice with the magnetic wallpaper group $G_{b}=p4gm\times Z_{2}^{\mathcal{T}}$. On the 2a Wyckoff positions A=(0,0) and B=(${1\over 2},{1\over 2}$) with site group $\mathscr{C}_{4}\times Z_{2}^{\mathcal{T}}$, we place spin-1/2 fermionic orbitals $[c_{\uparrow}^{\dagger},c_{\downarrow}^{\dagger}]$. Assuming SOC is negligible, then the spin-1/2 fermion forms a Kramers doublet under time reversal $\mathcal{T}$ and is invariant under the site group rotation $\mathscr{C}_{4}$. Due to the SU(2) spin rotation symmetry, the BdG Hamiltonian can be written in the reduced Nambu bases $\psi_{i}^{\dagger}=[c_{{A}\uparrow}^{\dagger},c_{{B}\uparrow}^{\dagger},c_{{A}\downarrow},c_{{B}\downarrow}]_{i},$ where the symmetry group $G_{f}$ is generated by $C^{\prime}_{4}=C_{4},\mathcal{T}^{\prime}=(IK||\mathcal{T}),\mathcal{M}_{x}^{\prime}=\{e^{-\mathrm{i}{\tau_{z}\over 2}\pi}||\mathcal{M}_{x}|{1\over 2},{1\over 2}\}$ (here $e^{-\mathrm{i}{\tau_{z}\over 2}\pi}$ is a charge operation defined in (8)) and translation, with
@@
and ${\widehat{\mathcal{C}}}^{2}=P_{f}$. Then we obtain the complete symmetry group $F^{c}_{f}$, whose band representation can be obtained according to App.C. Irreducible zero modes are found at $X=({1\over 2},0),Y=(0,{1\over 2}),M=({1\over 2},{1\over 2})$ points, where the $X$ and $Y$ points are related by $C^{\prime}_{4}$ symmetry.
where $\zeta=A,B$ and $\bar{\zeta}\neq\zeta$. The phase $e^{\mathrm{i}\theta_{i,j}}=\pm 1$ of the pairing term on the bond $(ij)$ alternatively equals to $1,-1$, as shown in Fig.1. After Fourier transformation, the Hamiltonian can be written in the bases $\psi_{k}^{\dagger}=[c_{kA\uparrow}^{\dagger},c_{kB\uparrow}^{\dagger},c_{-kA\downarrow},c_{-kB\downarrow}]$. The band structure with parameters $t=1,\Delta_{2}=1.5$ is shown in Figs.2(a) & 2(b), where the 2-fold degeneracy of the bands are due to the anti-unitary symmetry ${\mathcal{C}}\mathcal{I}\mathcal{T}^{\prime}$ with $\mathcal{I}={C^{\prime}}_{4}^{2}$, $({{\mathcal{C}}\mathcal{I}\mathcal{T}^{\prime}})^{2}=-1$. In the following, we analysis the protection of zero modes.
@@
$H_{\rm eff}=a\delta k_{1}\tau_{x}\otimes\nu_{y}+b\delta k_{2}\tau_{z}\otimes\nu_{y},$
where $a,b$ are constants, $k_{1}=k_{y},k_{2}=k_{x}$ for the X point and $k_{1}=k_{x},k_{2}=k_{y}$ for the Y point.
@@
$D(C^{\prime}_{4})=I_{2}\otimes\nu_{z},\ D(\mathcal{M}^{\prime}_{x})=\tau_{z}\otimes\nu_{y},\ D(\mathcal{T}^{\prime})K=I_{4}K,$
@@
$H_{\rm eff}=a(\delta k_{x}^{2}-\delta k_{y}^{2})\tau_{x}\otimes\nu_{x}+b\delta k_{x}\delta k_{y}\tau_{z}\otimes\nu_{x}.$
No zero modes appear at the $\Gamma$ point with $F_{f}^{c}(\Gamma)/Z_{2}^{f}=\mathscr{C}_{4v}\times Z_{2}^{\mathcal{C}}\times Z_{2}^{\mathcal{T}}\times Z_{2}^{\mathcal{P}}$. Compared with the M point, $C^{\prime}_{4}$ and $\mathcal{M}^{\prime}_{x}$ are represented differently as $D(C^{\prime}_{4})=I_{2}\otimes I_{2}$ and $D(\mathcal{M}^{\prime}_{x})=\mathrm{i}\tau_{z}\otimes\nu_{x}$, consequently the irrep at $\Gamma$ point correspond to irreducible nonzero modes.
Now we investigate the stability of the zero modes by adding symmetry breaking terms to the Hamiltonian. Interestingly, if one removes the mirror symmetry and preserves all the other symmetries by adding, for instance, the $\Delta_{0}$ perturbation term in Eq.(25), then the dispersion is qualitatively the same as Fig.2(a) and the 4-fold degenerate zero modes at the X, Y and M points still remain. However, unlike the case with $\Delta_{0}=0$ in which all of the 4-fold $\mathcal{C}$-minimal zero modes are irreducible, when $\Delta_{0}\neq 0$ the set of $\mathcal{C}$-minimal zero modes become a direct sum of two sets of 2-fold irreducible zero modes uncoupled to each other.
@@
Here the $\Delta_{0}$ term breaks the the mirror symmetry while the $\Delta_{1}$ term breaks both the mirror and $\mathcal{C}$ symmetries, and both terms preserve the $C^{\prime}_{4}$ and $\mathcal{T}^{\prime}$ symmetries. The Band structure of the Hamiltonian $H_{0}+H_{1}+H_{2}$ with $t=1,\Delta_{2}=1.5,\lambda=0.5,t_{0}=0.6,\Delta_{0}=2,\Delta_{1}=1.7$ are shown in Fig.3 with the cones protected by the combined $\mathcal{I}\mathcal{T}^{\prime}$ symmetry with $(\widehat{\mathcal{I}\mathcal{T}^{\prime}})^{2}=1$. The $\mathcal{I}\mathcal{T}^{\prime}$ symmetry gives rise to $\pi$-quantized Berry phase for loops in the BZ surrounding a single zero mode, which protects the cones from being solely gapped out. This mechanism makes the zero modes robust when the $\mathcal{I}\mathcal{T}^{\prime}$ symmetry is preserved.
@@
Finally, when the pairing terms are complex, then the time reversal symmetry $\mathcal{T}^{\prime}$ and $\mathcal{I}\mathcal{T}^{\prime}$ are broken and the zero modes disappear. The fully gapped bands becomes a topological SC/SF with nonzero Chern number. Due to the unbroken $C^{\prime}_{4}$ symmetry, the Chern number is either $4$ or $-4$. As shown in Fig.4, if we keep $\Delta_{1,2}$ to be real but vary the phase of $\Delta_{0}$, then the Chern number can be changed from 4 to $-4$ or vice versa. The phase transition occurs at $\theta_{0}={\pi}$ ($\Delta_{0}$ is real again), where the gap closes at 8 nodal points in the BZ.
From the above model, we can see that the zero modes are quit robust even in the presence of symmetry breaking perturbations. When the effective charge conjugation symmetry $\mathcal{C}$ is preserved, the zero modes appear at high symmetry points. When the $\mathcal{C}$ is violated, like in most realist SCs/SFs, the zero modes do not disappear at once as long as certain protecting symmetries are still preserved. The $\mathcal{C}$ symmetry is of theoretical significance in the study of BdG systems in the sense that it provides the physical origin of the zero modes when the $\mathcal{C}$ breaking terms are treated as perturbations. This provides a systematic way to investigate the physical properties of nodal SCs/SFs with gapless quasiparticle excitations.

### v3

@@ II.4 Conditions for the existence of Zero modes
Now we consider BdG systems on the lattice having a symmetry group $G_{f}$, and assume that a fixed number of complex fermion bases (namely local Wannier orbitals) are placed in certain Wyckoff positions in the unit cell. According to App.C, for any given momentum $\bm{k}$, the rep of $F_{f}(k)$ can be read out from the generalized band representation, and then we can know whether there exist zero modes at the $\bm{k}$. If zero modes do exist, the degeneracy must be even because the dimension of the Hamiltonian $\mathscr{H}_{k}$ is even and the nonzero energy levels appear in pairs $\{\varepsilon_{k},-\varepsilon_{k}\}$. In the following theorem 2 (see App.F for proof), however, we show that if the $G_{f}(k)$ is ‘diagonal’ in the particle-hole sector, then no stable zero modes can be found at a given momentum $\bm{k}$.
@@
For any BdG system whose symmetry group $G_{f}(k)$ is ‘diagonal’ in the particle-hole sector, all zero modes at the momentum point $\bm{k}$ are accidental and can be adiabatically lifted without breaking any symmetry.
However, situations will be changed if the system contains ‘non-diagonal’ charge symmetries like the (effective) charge conjugation symmetry. In App.G, it is proved that at a given $\tilde{\mathcal{P}}$ symmetric momentum point $\bm{k}$, one can always find (effective) charge-conjugation symmetries to turn a set of minimal nonzero modes into zero modes – the minimal ‘$\mathcal{C}$-zero modes’. If the $\mathcal{C}$-zero modes originate from irreducible nonzero modes of $F_{f}(k)$, then they form irreducible zero modes of the group $F^{c}_{f}(k)=F_{f}(k)+\mathcal{C}F_{f}(k)$; otherwise if the $\mathcal{C}$-zero modes come from RMNZM of $F_{f}(k)$, then they are called reducible minimal $\mathcal{C}$-zero modes. Then we have the following theorem 3.
@@
At a given $\tilde{\mathcal{P}}$-symmetric momentum $\boldsymbol{k}$ (i.e. $\tilde{\mathcal{P}}\in F_{f}(k)$) of an arbitrary BdG system, any minimal nonzero modes (except for one special case) located at the momentum $\bm{k}$ can be turned into zero modes by adding a single charge-conjugation symmetry to the original BdG system. The exception is when $\tilde{\mathcal{P}}$ is anti-unitary while $G_{f}(k)$ is unitary, and the irrep $D(F_{f}(k))$ carried by a set of irreducible nonzero modes is of $\mathbb{H}$ class, then one needs two charge-conjugation operations anti-commuting with each other to gain zero modes.
Charge conjugations are the simplest charge operations that are non-diagonal in the particle-hole sector. The above theorem can be generalized to cases where the charge conjugation operation $\mathcal{C}$ is replaced by general non-diagonal charge operations $\mathcal{D}$ in $SU(2)_{c}$, e.g. with $\mathcal{D}^{3}=E$ or $\mathcal{D}^{6}=E$.
On the other hand, zero modes can also appear at unfixed momenta. For instance, a level crossing of irreducible nonzero modes carrying two nonequivalent reps of $G_{f}(k)$ can give rise to zero modes on certain high symmetry line. So we have the following theorem:
@@
On a high symmetry line of the BZ having a unitary symmetry $\mathcal{M}$, if the eigenspaces $\{\varepsilon_{k},-\varepsilon_{k}\}$ of a minimal nonzero modes carry different quantum numbers (i.e. characters of irreps) of $\mathcal{M}$, then the level crossing of $\varepsilon_{k},-\varepsilon_{k}$ gives rise to a set of stable zero modes whose momentum is not fixed.
@@
(1) presence of one or two (effective) charge-conjugation symmetries giving rise to one or more sets of irreducible zero modes at a high symmetry point $\boldsymbol{k}$.
@@
(3) occurrence of $\pi$-quantized Berry phase around generic $\bm{k}$ point due to the ($\mathcal{I}^{\prime}\mathcal{T}^{\prime}$) symmetry with $(\mathcal{I}^{\prime}\mathcal{T}^{\prime})^{2}=1$. Here $\mathcal{I}^{\prime}$ and $\mathcal{T}^{\prime}$ respectively stand for the inversion and the time-reversal symmetries dressed by spin or/and charge operations.
The condition (1) can be verified using the band representation theory for SCs/SFs/QSLs provided in App.C. The last two conditions are illustrated via concrete lattice models in Sec.IV. Especially, the $\mathcal{I}^{\prime}\mathcal{T}^{\prime}$ symmetry enforced $\pi$-quantized Berry phase can give rise to nodal-point like quasiparticles in 2-dimensions and nodal-line structures in 3-dimensions. This is a topological origin of the zero modes (usually linear dispersive), and is closely related to the zero modes at generic $\bm{k}$ points enforced by symmetry indicators Slager et al. [2013], Kruthoff et al. [2017], Po et al. [2017], Bradlyn et al. [2017], Ono and Watanabe [2018], Geier et al. [2020], Ono et al. [2021], Ono and Shiozaki [2022]. The (projective) representation theory of symmetry groups discussed in the present work provides a different and complementary mechanism to obtain zero modes and gapless quasiparticles.
@@ III.1 $k\cdot p$ Hamiltonian
The above procedure can be applied to the 0-th order $k\cdot p$ terms, i.e., to judge if an irrep is simple or composite. This can be done by replacing the rep $({\rm v}\times{\rm{\overline{ph}}})$ in eq.(III.1) with the identity rep or the particle-hole rep. If $a_{I}=1,a_{\rm ph}=0$, then the irrep is simple and cannot form nonzero modes by itself; if $a_{I}=1,a_{\rm ph}\neq 0$, then the irrep is composite which corresponds to a set of irreducible nonzero modes.
@@
Table 2: Linear Reps of the lattice rotation ($l$ stands for angular momentum), time reversal $\mathcal{T}^{\prime}$, spatial inversion $\mathcal{I}^{\prime}$, charge conjugation $\mathcal{C}$ and particle-hole transformation $\mathcal{P}$ carried by probe fields: chemical potential $\lambda$, momentum $\boldsymbol{k}$, gradience of temperature $\boldsymbol{\nabla}T$, magnetic field $\boldsymbol{B}$, quadratic dispersion $k_{\alpha}k_{\beta}$, strain field $\epsilon_{\alpha\beta}$, and so on. The last column is the ‘particle-hole rep’ of operations in $F_{f}(k)$.
prob fields $\lambda$ $\boldsymbol{\nabla}T$ $\boldsymbol{B}$ $\boldsymbol{k}$ $k_{\alpha}k_{\beta}$ $\epsilon_{\alpha\beta}$ $D^{\rm{(ph)}}$
@@ III.2 Physical properties
The above method can also be applied to study the physical properties of the quasiparticles, namely, the response to external fields $\boldsymbol{f}$, such as the gradience of temperature $\boldsymbol{\nabla}T$, magnetic field $\boldsymbol{B}$, strain field $\epsilon_{\alpha\beta}$, and so on. As summarized in Tab.2, the prob fields carry linear reps of $g\in F_{f}(k)$, namely
@@
Similar to (18), one has $D(g)K_{s(g)}\Lambda_{\alpha}K_{s(g)}D(g)^{-1}=\sum_{\beta}D^{{\rm(ph)}\times(\mu)}_{\beta\alpha}(g)\Lambda_{\beta}$. The matrices $\Lambda_{\alpha}$ can be obtained using the method introduced in the previous subsection. Notice that electric fields $\boldsymbol{E}$ are not an applicable probe fields, since electric fields are screened in SCs and do not couple to the charge neutral SFs/QSLs. Instead, temperature gradience $\boldsymbol{\nabla}T$ is an ideal probe field for SCs/SFs/QSLs, and $\boldsymbol{\nabla}T$ varies under symmetry operations in the same way as $\boldsymbol{E}$. Magnetic fields can be applied as Zeeman fields coupling to SCs/SFs/QSLs. Despite the Meisner effect of SCs, magnetic fields can be applied along in-plane directions for thin films of SCs.
Due to the ‘particle-hole symmetry’ $\mathcal{P}$ and the effective charge conjugation symmetry $\mathcal{C}$, the response of the Bogoliubov quasiparticles to external field is different from quasiparticles in semimetals. For instance, the chemical potential $\lambda$ is a scaler field under space-time operations, but it does not carry an identity rep of the group $F_{f}^{c}(k)=F_{f}(k)+\mathcal{C}F_{f}(k)$, hence the response matrix $\Lambda_{0}$ is not simply an identity matrix. Consequently changing the chemical potential may either shift the position of the nodal point or fully gap out the quasiparticle.
Generally, the external fields may give mass to the gapless quasiparticles. The resulting gapped state may have nontrivial topological properties, such as nontrivial Chern numbers. These information are maintained in the $k\cdot p$ model and the response matrix mentioned above. In the following section, we will illustrate the above results by concrete lattice models.
@@ IV Lattice Models: gapless quasiparticles in BdG systems
In this section, we present four models in two-dimensional translationally symmetric lattice systems to illustrate the mechanisms and methods proposed in the previous sections. The first three models discuss zero modes protected by an effective charge conjugation $\mathcal{C}$, while the fourth model is a mean field Hamiltonian of $Z_{2}$ QSLs in which the projective symmetry group contains symmetry elements that are non-diagonal in the particle-hole sector.
@@ IV.1 Square lattice: $G_{b}=p4gm\times Z_{2}^{\mathcal{T}}$
The first model is a spin-1/2 SC/SF of the CI class. We consider the square lattice with the magnetic wallpaper group $G_{b}=p4gm\times Z_{2}^{\mathcal{T}}$. On the 2a Wyckoff positions A=(0,0) and B=(${1\over 2},{1\over 2}$) with site group $\mathscr{C}_{4}\times Z_{2}^{\mathcal{T}}$, we place spin-1/2 fermionic orbitals $[c_{\uparrow}^{\dagger},c_{\downarrow}^{\dagger}]$. Assuming SOC is negligible, then the spin-1/2 fermion forms a Kramers doublet under time reversal $\mathcal{T}$ and is invariant under the site group rotation $\mathscr{C}_{4}$. Due to the SU(2) spin rotation symmetry, the BdG Hamiltonian can be written in the reduced Nambu bases $\psi_{i}^{\dagger}=[c_{{A}\uparrow}^{\dagger},c_{{B}\uparrow}^{\dagger},c_{{A}\downarrow},c_{{B}\downarrow}]_{i},$ where the symmetry group $G_{f}$ is generated by $C_{4}=C_{4},\mathcal{T}^{\prime}=(IK||\mathcal{T}),\mathcal{M}_{x}^{\prime}=\{e^{-\mathrm{i}{\tau_{z}\over 2}\pi}||\mathcal{M}_{x}|{1\over 2},{1\over 2}\}$ (here $e^{-\mathrm{i}{\tau_{z}\over 2}\pi}$ is a charge operation defined in (8)) and translation, with
@@
and ${\widehat{\mathcal{C}}}^{2}=P_{f}$. Then we obtain the complete symmetry group $F^{c}_{f}$, whose band representation can be obtained according to App.C. Irreducible zero modes are found at $X=({1\over 2},0),Y=(0,{1\over 2}),M=({1\over 2},{1\over 2})$ points, where the $X$ and $Y$ points are related by $C_{4}$ symmetry.
@@
where $\zeta=A,B$ and $\bar{\zeta}\neq\zeta$. The phase $e^{\mathrm{i}\theta_{i,j}}=\pm 1$ of the pairing term on the bond $(ij)$ alternatively equals to $1,-1$, as shown in Fig.1. After Fourier transformation, the Hamiltonian can be written in the bases $\psi_{k}^{\dagger}=[c_{kA\uparrow}^{\dagger},c_{kB\uparrow}^{\dagger},c_{-kA\downarrow},c_{-kB\downarrow}]$. The band structure with parameters $t=1,\Delta_{2}=1.5$ is shown in Figs.2(a) & 2(b), where the 2-fold degeneracy of the bands are due to the anti-unitary symmetry ${\mathcal{C}}\mathcal{I}\mathcal{T}^{\prime}$ with $\mathcal{I}={C}_{4}^{2}$, $({{\mathcal{C}}\mathcal{I}\mathcal{T}^{\prime}})^{2}=-1$. In the following, we analysis the protection of zero modes.
@@
$H_{\rm eff1}=a\delta k_{1}\tau_{x}\otimes\nu_{y}+b\delta k_{2}\tau_{z}\otimes\nu_{y},$
where $a,b$ are constants, $k_{1}=k_{y},k_{2}=k_{x}$ for the X point and $k_{1}=k_{x},k_{2}=k_{y}$ for the Y point. Furthermore, the response to the external fields can also be derived based on the theory discussed in Sec.III. For instance, the response to the temperature gradient $\boldsymbol{\nabla}T$ at the $Y$ point is linear, which is given by
$H_{\boldsymbol{\nabla}T}=a^{\prime}\partial_{x}T\tau_{z}\otimes\nu_{x}+b^{\prime}\partial_{y}T\tau_{x}\otimes\nu_{x}.$
@@
$D(C_{4})=I_{2}\otimes\nu_{z},\ D(\mathcal{M}^{\prime}_{x})=\tau_{z}\otimes\nu_{y},\ D(\mathcal{T}^{\prime})K=I_{4}K,$
@@
$H_{\rm eff2}=c(\delta k_{x}^{2}-\delta k_{y}^{2})\tau_{x}\otimes\nu_{x}+d\delta k_{x}\delta k_{y}\tau_{z}\otimes\nu_{x}.$
The lowest order response to the temperature gradient $\boldsymbol{\nabla}T$ at the $M$ point is of second order, which is given by
$H_{\boldsymbol{\nabla}T}=c^{\prime}({\partial_{x}T}^{2}-{\partial_{y}T}^{2})\tau_{x}\otimes\nu_{x}+d^{\prime}{\partial_{x}T}{\partial_{y}}T\tau_{z}\otimes\nu_{x}.$
No zero modes appear at the $\Gamma$ point with $F_{f}^{c}(\Gamma)/Z_{2}^{f}=\mathscr{C}_{4v}\times Z_{2}^{\mathcal{C}}\times Z_{2}^{\mathcal{T}}\times Z_{2}^{\mathcal{P}}$. Compared with the M point, $C_{4}$ and $\mathcal{M}^{\prime}_{x}$ are represented differently as $D(C_{4})=I_{2}\otimes I_{2}$ and $D(\mathcal{M}^{\prime}_{x})=\mathrm{i}\tau_{z}\otimes\nu_{x}$, consequently the irrep at $\Gamma$ point correspond to irreducible nonzero modes.
Now we investigate the stability of the zero modes by adding symmetry breaking terms to the Hamiltonian. Interestingly, if one removes the mirror symmetry and preserves all the other symmetries by adding, for instance, the $\Delta_{0}$ perturbation term in Eq.(25), then the dispersion is qualitatively the same as Fig.2(a) and the 4-fold degenerate zero modes at the X, Y and M points still remain. However, unlike the case $\Delta_{0}=0$ where the set of 4-fold $\mathcal{C}$-zero modes is irreducible, when $\Delta_{0}\neq 0$ the 4-fold zero modes are reducible minimal $\mathcal{C}$-zero modes.
@@
Here the $\Delta_{0}$ term breaks the the mirror symmetry while the $\Delta_{1}$ term breaks both the mirror and $\mathcal{C}$ symmetries, and both terms preserve the $C_{4}$ and $\mathcal{T}^{\prime}$ symmetries. The Band structure of the Hamiltonian $H_{0}+H_{1}+H_{2}$ with $t=1,\Delta_{2}=1.5,\lambda=0.5,t_{0}=0.6,\Delta_{0}=2,\Delta_{1}=1.7$ are shown in Fig.3 with the cones protected by the combined $\mathcal{I}\mathcal{T}^{\prime}$ symmetry with $(\widehat{\mathcal{I}\mathcal{T}^{\prime}})^{2}=1$. The $\mathcal{I}\mathcal{T}^{\prime}$ symmetry gives rise to $\pi$-quantized Berry phase for loops in the BZ surrounding a single zero mode, which protects the cones from being solely gapped out. This mechanism makes the zero modes robust when the $\mathcal{I}\mathcal{T}^{\prime}$ symmetry is preserved.
@@
Finally, when the pairing terms are complex, then the time reversal symmetry $\mathcal{T}^{\prime}$ and $\mathcal{I}\mathcal{T}^{\prime}$ are broken and the zero modes disappear. The fully gapped bands becomes a topological SC/SF with nonzero Chern number. Due to the unbroken $C_{4}$ symmetry, the Chern number is either $4$ or $-4$. As shown in Fig.4, if we keep $\Delta_{1,2}$ to be real but vary the phase of $\Delta_{0}$, then the Chern number can be changed from 4 to $-4$ or vice versa. The phase transition occurs at $\theta_{0}={\pi}$ ($\Delta_{0}$ is real again), where the gap closes at 8 nodal points in the BZ.
From the above model, we can see that the zero modes are quit robust even in the presence of symmetry breaking perturbations. When the effective charge conjugation symmetry $\mathcal{C}$ is preserved, the zero modes appear at high symmetry points. When the $\mathcal{C}$ is violated, like in most realist SCs/SFs, the zero modes do not disappear at once as long as certain protecting symmetries are still preserved. The $\mathcal{C}$ symmetry is of theoretical significance for BdG systems in the sense that it provides the physical origin of the zero modes when the $\mathcal{C}$ breaking terms are treated as perturbations. This provides a systematic way to investigate the physical properties of nodal SCs/SFs with gapless quasiparticle excitations.